# Classification of superpotentials for cohomogeneity one Ricci solitons

### v1

@@ Abstract.
We classify superpotentials for the Hamiltonian system corresponding to the cohomogeneity one gradient Ricci soliton equations. Aside from recovering known examples of superpotentials for steady solitons, we find a new superpotential on a specific case of the Bérard-Bergery–Calabi ansatz. The latter is used to obtain an explicit formula for a steady complete soliton with an equidistant family of hypersurfaces given by circle bundles over $S^{2}\times S^{2}$. There are no superpotentials in the non-steady case in dimensions greater than 2, even if polynomial coefficients are allowed. We also briefly discuss generalised first integrals and the limitations of some known methods of finding them.
The author is supported by scholarships from the Engineering and Physical Sciences Research Council (EPSRC) and the Fonds de recherche du Québec – Nature et technologies (FRQNT)
@@ 1. Introduction
The Einstein equations can be significantly simplified when the underlying manifold is assumed to have certain symmetries. For instance, on a homogeneous space $G/K$, all curvature quantities can be computed from the Lie algebra $\mathfrak{g}$ of $G$ and the critical points of the scalar curvature functional can be used to find Einstein manifolds [10]. A natural generalisation of homogeneous spaces is the notion of a cohomogeneity one space, i.e. $G$ acts on the manifold via isometries with principal orbits which are hypersurfaces. In this context, the Einstein equations simplify to a system of ordinary differential equations (ODEs) in the transverse direction to the principal orbits [1, 8]. In certain cases, the abovementioned ODEs are explicitly integrable, producing new examples of Einstein manifolds [1].
@@
When $\epsilon<0$, $\epsilon=0$ and $\epsilon>0$, the soliton is said to be shrinking, steady and expanding respectively. Ricci solitons generate simple solutions to the Ricci flow. They are a generalisation of Einstein metrics, as (1.1) reduces to the Einstein condition when $X$ is a Killing vector field of $(M,\hat{g})$. In the cohomogeneity one context, (1.1) also simplifies to a system of ODEs [7] and a Hamiltonian approach is also possible [2], with the configuration space extended to include the soliton potential. Generalised first integrals and superpotentials are also defined in [2], and some examples are given, leading to explicit formulae for e.g. the 5-dimensional Bryant soliton. The main objective of the present work is to classify superpotentials for cohomogeneity one solitons using constraints from convex geometry, analogously to the work done in [4, 5, 6] for the Ricci-flat condition.
Our main result is Theorem 3.17, which states that the only scalar curvature weight vector configurations admitting a superpotential are the Bérard-Bergery–Calabi ansatz and a small number of low-dimensional examples. The result is analogous to Theorems 1.12 and 1.14 of [6] in the Ricci-flat case. We also consider the more restrictive non-steady case, which admits no superpotentials with constant coefficients, and just one, with 1-dimensional fibres, when we allow the coefficients to be polynomials.
All superpotentials in the classification had previously been found in [2], but one. We integrate the first-order subsystem corresponding to the latter to obtain an explicit expression for a 6-dimensional Ricci soliton, and verify that suitable boundary conditions at the singular orbit are satisfied to ensure smoothness of the metric. There is also a brief discussion of generalised first integrals. Thus, the present work is to a certain extent a sequel of [2].
The paper is organised as follows. First, we introduce in Section 2 the Hamiltonian formalism in question. Section 3, composing the bulk of the paper, details the classification of superpotentials in the steady case. In Section 4, we discuss the non-steady case and superpotentials with polynomial coefficients. In Section 5, we integrate the first-order subsystem for the newly found superpotential and write down explicit formulae for the soliton. Finally, in Section 6, we discuss the limitations of strategies used in [2] to find generalised first integrals.
@@ 2. Hamiltonian formalism for cohomogeneity one gradient solitons
A gradient Ricci soliton on a smooth manifold $M$ is a pair $(\hat{g},u)$ consisting of a Riemannian metric $\hat{g}$ and a real-valued function $u$ on $M$ such that for some $\epsilon\in\mathbb{R}$,
@@
Let $G$ be a compact Lie group acting on $(M,\hat{g})$ via isometries with cohomogeneity one. There is a dense open submanifold $M_{0}\subset M$ diffeomorphic to $I\times G/K$, where $K\subset G$ is a closed subgroup and $I\subset\mathbb{R}$ is an interval with coordinate $t$. Each $\{t\}\times G/K$ is a principal orbit, and we choose a unit speed geodesic $\gamma(t)$ intersecting each principal orbit orthogonally. Then a general cohomogeneity one metric on $M_{0}$ takes the form
@@
(2.3) $ric-\operatorname{tr}(L)L-\dot{L}+\dot{u}L+\frac{\epsilon}{2}I_{n}$ $=0$
@@
Given a principal orbit $G/K$, fix an $\mathrm{Ad}(K)$-invariant decomposition
$\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$
of the Lie algebra $\mathfrak{g}$ of $G$, so that $\mathfrak{p}$ is identified with the tangent space to a principal orbit. We define a Hamiltonian formalism for the system with configuration space $S^{2}_{+}(\mathfrak{p})^{K}\times\mathbb{R}$, where the first factor is identified via a fixed metric $Q$ on $\mathfrak{p}$ to the space of $G$-invariant Riemannian metrics on $G/K$, and the second factor is the soliton potential. In the monotypic case, we let
@@ 3. Superpotentials for steady solitons
(3.2) $f=\sum_{\mathbf{c}\in\mathcal{C}}f_{\mathbf{c}}e^{\mathbf{c}\cdot\mathbf{q}},\quad\mathrm{with}\quad|\mathcal{C}|<\infty.$
Then (3.1) becomes
@@ 3.1. Classification of superpotentials
Several examples of superpotentials have been found in [2]. In this section, we obtain a classification result, Theorem 3.17, for superpotentials of the form (3.2) under mild hypotheses on the homogeneous space $G/K$.
The superpotential condition is much more rigid for the Ricci soliton equations than for the Ricci-flat equations for a few reasons. First, Proposition 3.3 states that superpotentials for solitons must have a null vector, a result analogous to the Einstein condition with a cosmological constant [4, Theorem 10.1]. Also, the presence of the “energy” term $E$ in the extended scalar curvature formula, together with Lemma 3.2(i), implies that superpotentials for solitons must contain vectors which extend beyond the hyperplane of vectors whose components sum to $-1$, which contains all scalar curvature vectors. This additional dimension, together with the natural hypothesis that $\dim\mathrm{conv}(\mathcal{W})=r-1$, make (3.3) considerably more restrictive than in the Ricci-flat case and are key to obtaining a full classification with significantly less effort than was necessary in [4, 5, 6].
@@
which is the equation of a slanted cone in $\mathbb{R}^{r+1}$. We define the affine hyperplane $P=\{\mathbf{v}\in\mathbb{R}^{r+1}\>|\>v_{r+1}=-1\}=\{\frac{1}{2}(\mathbf{d}+\mathbf{x})\in\mathbb{R}^{r+1}\>|\>x_{r+1}=0\}$. Our first objective is to show that $\mathcal{C}\subset P$.
@@
From this point on, we will often write “vertex” to mean “vertex of $\mathrm{conv}(\mathcal{C})$” and “null” to mean “$J$-null”. We will assume that all weight vectors are nontrivial, i.e. $\textbf{0}\notin\mathcal{C}$.
@@
Suppose there are 3 or more. Since each of them is connected to every other one by edges, Lemma 3.5 implies that they are all proportional and thus collinear, which is impossible. ∎
@@
    Type I: one entry of w is $-1$, the others are zero. These are denoted $(-1^{i})$ for some $i=1,\dots,r$.
@@
In our classification of superpotentials, we will assume that $\mathrm{conv}(\mathcal{W})\subset P$ has nonzero $(r-1)$-dimensional measure. This is the same hypothesis as in the main theorem of [5]. It holds, for instance, if $G$ is semisimple. In the case that the isotropy group $K$ is connected, this is not so restrictive, due to the following result.
@@
We begin by noting that there is at least one type I vector, say $(-1^{r})\in\mathcal{W}$ up to permutation of indices. Let $\textbf{w}_{1},\dots,\textbf{w}_{n}$ be the other vectors of $\mathcal{W}$. First, we note that it suffices to show that the image of $\mathrm{conv}(\mathcal{W})$ under the affine transformation which sends the last ($r$-th) coordinate to zero has nonzero measure as a subset of $\{x_{r}=0\}\cong\mathbb{R}^{r-1}$. Since $(-1^{r})$ is sent to 0 by this transformation, it suffices to show that the matrix whose rows are the components of $\textbf{w}_{1},\dots,\textbf{w}_{n}$ has rank $r-1$.
@@
Assume for sake of contradiction that there are vertices on only one side of $P$, then there are either one or two. Suppose there is just one, call it c. Then it is adjacent to every vertex of $\mathrm{conv}(\mathcal{C})$ in $P$, which span an $r$-dimensional subset of $P$. However, they must all lie in $\mathrm{ker}J(\textbf{c},\cdot)\cap P\subsetneq P$, which is a contradiction. To see this, use the closest point argument on the edge connecting each such vertex with c, and use the fact that c is null. Suppose that there are two, then by Lemma 3.5 they are given by c and $\lambda\textbf{c}$ for some $\lambda\in\mathbb{R}$. Then each of the vertices in $P$ must be adjacent to at least one of these, and thus be in $\mathrm{ker}J(\textbf{c},\cdot)\cap P\subsetneq P$, a contradiction. ∎
@@
There is no null vertex in $P$, so since $\mathrm{conv}(\frac{1}{2}(\textbf{d}+\tilde{\mathcal{W}}))\subset\mathrm{conv}(\mathcal{C})\cap P$ and every vertex outside of $\frac{1}{2}(\textbf{d}+\tilde{\mathcal{W}})$ is null by Lemma 3.2, we have that $\mathrm{conv}(\frac{1}{2}(\textbf{d}+\tilde{\mathcal{W}}))=\mathrm{conv}(\mathcal{C})\cap P=\mathrm{conv}(\mathcal{C}\cap P)$, and that every vertex of $\mathrm{conv}(\frac{1}{2}(\textbf{d}+\tilde{\mathcal{W}}))$ is a vertex of $\mathcal{C}$. From this point on, we may follow the proof of Lemma 5.7 in [2], which we reproduce here for completeness. Since 0 is a vertex of $\mathrm{conv}(\tilde{\mathcal{W}})$ and there are no null vertices in $P$, there is a $\textbf{w}\in\mathcal{W}$ such that the vertices $\textbf{c}_{0}=\frac{1}{2}\textbf{d}$ and $\textbf{c}_{1}=\frac{1}{2}(\textbf{d}+\textbf{w})$ form an edge of $\mathrm{conv}(\mathcal{C})$. Then $J(\textbf{c}_{0},\textbf{c}_{0})=\frac{1}{4}$ and $J(\textbf{c}_{0},\textbf{c}_{1})=\frac{1}{4}$, and thus $J(\textbf{c}_{0},\cdot)$ is constant and nonzero on the line segment $\overline{\textbf{c}_{0}\textbf{c}_{1}}$. Consider the element of $\mathcal{C}$ on $\overline{\textbf{c}_{0}\textbf{c}_{1}}$ which is closest to $\textbf{c}_{0}$. It satisfies the unique sum condition with $\textbf{c}_{0}$, which contradicts Lemma 3.2.
It only remains to check the case where there exists a pair of vertices, a and c, one on each side of $P$, such that $\overline{\textbf{a}\textbf{c}}$ is an edge of $\mathrm{conv}(\mathcal{C})$. Then either $\textbf{a}+\textbf{c}\in\textbf{d}+\tilde{\mathcal{W}}$ or a and c are proportional, as can be seen as follows.
@@
If $\textbf{x}_{a}\neq\textbf{c}$, we note that $\textbf{a},\textbf{x}_{a}$ satisfy the unique sum condition. Similarly, we may define $\textbf{x}_{c}\in\overline{\textbf{a}\textbf{c}}$ such that $\textbf{c},\textbf{x}_{c}$ satisfy the unique sum condition. Since a and c lie on different sides of $P$, we have that at least one of $\textbf{a}+\textbf{x}_{a}$ or $\textbf{c}+\textbf{x}_{c}$ is not in $\textbf{d}+\tilde{\mathcal{W}}\subset P$; assume without loss of generality that it is $\textbf{a}+\textbf{x}_{a}$. Then, by Lemma 3.2, $J(\textbf{a},\textbf{x}_{a})=0$, and by linearity $J(\textbf{a},\textbf{c})=0$, implying that a and c are proportional.
@@
Let $\textbf{f}=\frac{1}{2}(\textbf{d}+\textbf{w})$ be a non-null vertex which shares an edge with both a and c. Let $\textbf{x}_{a}$ denote the point of $\mathcal{C}$ on $\overline{\textbf{a}\textbf{f}}$ closest to a, and let $\textbf{x}_{c}$ denote the point of $\mathcal{C}$ on $\overline{\textbf{c}\textbf{f}}$ which is closest to c. Then we must have by Lemma 3.2 that $J(\textbf{a},\textbf{x}_{a})=0$ and $J(\textbf{c},\textbf{x}_{c})=0$. Since $J(\textbf{a},\cdot)$ and $J(\textbf{c},\cdot)$ are affine functions on the abovementioned edges and both a and c are null vertices, we must have $J(\textbf{a},\textbf{f})=J(\textbf{c},\textbf{f})=0$. This is impossible, as $J(\textbf{a},\textbf{f})+J(\textbf{c},\textbf{f})=J(\textbf{d},\frac{1}{2}(\textbf{d}+\textbf{w}))=\frac{1}{2}$. Therefore $\textbf{a}+\textbf{c}\notin\textbf{d}+\tilde{\mathcal{W}}$.
In the final case, namely that a and c are proportional, then by Lemma 3.9 there is only one vertex on each side of $P$. Furthermore, $\frac{\textbf{d}}{2}$ is not a vertex, since otherwise there must be a non-null vertex sharing an edge with $\frac{1}{2}\textbf{d}$, which is impossible.
Let =$\ell=\{\lambda\textbf{a}:\lambda\in\mathbb{R}\}$ be the (null) line passing through a and c, intersecting $P$ at $\mathbf{l}$. Then $\textbf{l}\in\{s\geq 0\}$. Since all vertices of $\mathrm{conv}(\mathcal{C})\cap P$ except possibly l are in $\{s=-1\}$, but $\frac{1}{2}\textbf{d}\in\mathrm{conv}(\mathcal{C})\cap P$, we have that $\textbf{l}\in\{s>0\}$.
Note that $\mathrm{conv}(\mathcal{C})$ has nonzero $(r+1)$-dimensional measure, since it contains the line $\overline{\textbf{a}\textbf{c}}$ through the origin as well as $\mathrm{conv}(\tilde{\mathcal{W}})$, which is an $r$-dimensional subset of $P$. Since all vertices of $\mathrm{conv}(\mathcal{C})$ other than a and c are in $\{s=-1\}\cap P$, which is $(r-1)$-dimensional, there is a set $\mathcal{V}=\{\textbf{b}_{i}\}_{i=1}^{r}\subset\{s=-1\}\cap P$ of vertices adjacent to a such that $\dim\mathrm{conv}(\mathcal{V})=r-1$. Applying the closest point argument, we get $J(\textbf{a},\textbf{b}_{i})=0$ for each $i$. The $(r-1)$-dimensional affine subspace $\ker J(\textbf{a},\cdot)\cap P$ contains both $\mathcal{V}$ and $\textbf{l}\in\overline{\textbf{a}\textbf{c}}\cap P$, but $\textbf{l}\notin\{s=-1\}$, which is a contradiction.
@@
There cannot be a non-null vertex adjacent to $\frac{1}{2}\textbf{d}$. Let a be a vertex adjacent to $\frac{1}{2}\textbf{d}$, then $J(\textbf{a},\textbf{a})=0$ and $J(\frac{1}{2}\textbf{d},\textbf{a})=\frac{1}{4}$, so that by the closest point argument and Lemma 3.2 we must have that the line segment between them is empty and that $\textbf{a}=\frac{1}{2}\textbf{d}+\textbf{w}$ for some $\textbf{w}\in\mathcal{W}$. All such vertices $\textbf{a}_{i}$, of which in all nontrivial cases $r\geq 2$ there exists more than one, therefore lie in $\{s=-2\}$. Consequently, there are at least two vertices in the half-plane $\{s<-1\}$, which are connected by edges lying entirely in $\{s<-1\}$. Thus, there is least one pair of adjacent (null) vertices in $\{s<-1\}$, which is impossible by the closest point argument and Lemma 3.2. ∎
@@
If there are 2 vertices in $\{s<-1\}$, then by the same argument they must sum to an element of $\textbf{d}+\tilde{\mathcal{W}}$ in $\{s<-1\}$, which is impossible. Therefore $b\leq 1$.
@@
More explicitly, we are saying that all vertices of $\mathrm{conv}(\mathcal{W})=\mathrm{conv}(\mathcal{C})\cap\{s=-1\}$ must take the form $\frac{1}{2}(\textbf{d}+\textbf{x})$, where

### v2

@@ Abstract.
We classify superpotentials for the Hamiltonian system corresponding to the cohomogeneity one gradient Ricci soliton equations. Aside from recovering known examples of superpotentials for steady solitons, we find a new superpotential on a specific case of the Bérard Bergery–Calabi ansatz. The latter is used to obtain an explicit formula for a steady complete soliton with an equidistant family of hypersurfaces given by circle bundles over $S^{2}\times S^{2}$. There are no superpotentials in the non-steady case in dimensions greater than 2, even if polynomial coefficients are allowed. We also briefly discuss generalised first integrals and the limitations of some known methods of finding them.
The author is supported by the Engineering and Physical Sciences Research Council [grant number EP/W524311/1] and the Fonds de recherche du Québec – Nature et technologies [numéro de dossier 331732].
Keywords: Cohomogeneity one, Ricci solitons, Hamiltonian systems, superpotentials
MSC codes: 53C25, 53C30, 37J35
@@ 1. Introduction
The Einstein equations can be significantly simplified when the underlying manifold is assumed to have certain symmetries. For instance, on a homogeneous space $G/K$, all curvature quantities can be computed from the Lie algebra $\mathfrak{g}$ of $G$ and the critical points of the scalar curvature functional can be used to find Einstein manifolds [10]. A natural generalisation of homogeneous spaces is the notion of a cohomogeneity one space, i.e. $G$ acts on the manifold via isometries with codimension one principal orbits. In this context, the Einstein equations simplify to a system of ordinary differential equations (ODEs) in the transverse direction to the principal orbits [1, 8]. In certain cases, the abovementioned ODEs are explicitly integrable, producing new examples of Einstein manifolds [1].
@@
When $\epsilon<0$, $\epsilon=0$ or $\epsilon>0$, the soliton is said to be shrinking, steady or expanding respectively. Ricci solitons generate simple solutions to the Ricci flow. They are a generalisation of Einstein metrics, as (1.1) reduces to the Einstein condition when $X$ is a Killing vector field of $(M,\hat{g})$. In the cohomogeneity one context, (1.1) also simplifies to a system of ODEs [7] and a Hamiltonian approach is also possible [2], with the configuration space extended to include the soliton potential. Generalised first integrals and superpotentials are also defined in [2], and some examples are given, leading to explicit formulae for e.g. the 5-dimensional Bryant soliton. The main objective of the present work is to classify superpotentials for cohomogeneity one solitons using constraints from convex geometry, analogously to the work done in [4, 5, 6] for the Ricci-flat condition.
Our main result is Theorem 3.17, which states that the only scalar curvature weight vector configurations admitting a superpotential are the Bérard Bergery–Calabi ansatz and a small number of low-dimensional examples. The result is analogous to Theorems 1.12 and 1.14 of [6], which classify superpotentials for the cohomogeneity one Ricci-flat Einstein equations. We also consider the more restrictive non-steady case, which admits no superpotentials with constant coefficients, and just one, with 1-dimensional fibres, when we allow the coefficients to be polynomials.
All superpotentials in the classification had previously been found in [2], except the one given in Theorem 3.17(5). We integrate the first-order subsystem corresponding to the latter and verify that suitable boundary conditions at the singular orbit are satisfied to ensure smooth extension of the metric. As a result, we obtain explicit formulae for a steady complete soliton on the space foliated by hypersurfaces given by $U(1)$-bundles over $S^{2}\times S^{2}$, detailed in Proposition 5.1. There is also a brief discussion of generalised first integrals. Thus, the present work is to a certain extent a sequel of [2].
The paper is organised as follows. First, we introduce in Section 2 the Hamiltonian formalism in question. Section 3, composing the bulk of the paper, details the classification of superpotentials in the steady case. In Section 4, we discuss the non-steady case and superpotentials with polynomial coefficients. In Section 5, we integrate the first-order subsystem for the newly found superpotential and write down explicit formulae for the corresponding $6$-dimensional soliton. Finally, in Section 6, we discuss the limitations of strategies used in [2] to find generalised first integrals.
@@ 2. Hamiltonian formalism for cohomogeneity one gradient solitons
A gradient Ricci soliton on a smooth manifold $M$ consists of a Riemannian metric $\hat{g}$ and a real-valued function $u$ on $M$ such that for some $\epsilon\in\mathbb{R}$,
@@
The vector field $X$ in (1.1) is given by the gradient of $u$.
Let $G$ be a compact Lie group acting on $(M,\hat{g})$ via isometries with cohomogeneity one. There is a dense open submanifold $M_{0}\subset M$ diffeomorphic to $I\times G/K$, where $K\subset G$ is a closed subgroup and $I\subset\mathbb{R}$ is an interval with coordinate $t$. Each $\{t\}\times G/K$ is a principal orbit, and we choose a unit speed geodesic $\gamma(t)$ intersecting each principal orbit $\hat{g}$-orthogonally. Then, a general cohomogeneity one metric on $M_{0}$ takes the form
@@
(2.3) $ric-\operatorname{tr}(L)L-\dot{L}+\dot{u}L+\frac{\epsilon}{2}I_{n}$ $=0,$
@@
Given a principal orbit $G/K$, let $\mathfrak{g}$ and $\mathfrak{k}$ denote the Lie algebras of $G$ and $K$ respectively. Fix an $\mathrm{Ad}(K)$-invariant decomposition
$\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p},$
so that $\mathfrak{p}$ is identified with the tangent space to a principal orbit. We define a Hamiltonian formalism for the system with configuration space $S^{2}_{+}(\mathfrak{p})^{K}\times\mathbb{R}$, where the first factor is identified via a fixed metric $Q$ on $\mathfrak{p}$ to the space of $G$-invariant Riemannian metrics on $G/K$, and the second factor is the soliton potential. In the monotypic case, we let
@@ 3. Superpotentials for steady solitons
(3.2) $f=\sum_{\mathbf{c}\in\mathcal{C}}f_{\mathbf{c}}e^{\mathbf{c}\cdot\mathbf{q}}\quad\mathrm{with}\quad|\mathcal{C}|<\infty,$
for some constants $f_{\textbf{c}}\neq 0$. Then (3.1) becomes
@@ 3.1. Classification of superpotentials
Several examples of superpotentials have been found in [2]. In this subsection, we obtain a classification result, Theorem 3.17, for superpotentials of the form (3.2) under mild hypotheses on the homogeneous space $G/K$.
The superpotential condition is much more rigid for the Ricci soliton equations than for the Ricci-flat equations for a few reasons. First, Proposition 3.3 states that superpotentials for (non-Ricci-flat) solitons must have a null vector, a result analogous to the case of the Einstein condition with a cosmological constant [4, Theorem 10.1]. Also, the presence of the “energy” term $E$ in the extended scalar curvature formula, together with Lemma 3.2(i), implies that superpotentials for solitons must contain vectors which extend beyond the hyperplane of vectors whose components sum to $-1$, which contains all scalar curvature vectors. This additional dimension, together with the natural hypothesis that $\dim\mathrm{conv}(\mathcal{W})=r-1$, make (3.3) considerably more restrictive than in the Ricci-flat case and are key to obtaining a full classification with significantly less effort than was necessary in [4, 5, 6].
@@
which is the equation of a slanted cone in $\mathbb{R}^{r+1}$. We define the affine hyperplane
$P=\{\mathbf{v}\in\mathbb{R}^{r+1}\>|\>v_{r+1}=-1\}=\left\{\frac{1}{2}(\mathbf{d}+\mathbf{x})\in\mathbb{R}^{r+1}\>\Big{|}\>x_{r+1}=0\right\}.$
Our first objective is to show that $\mathcal{C}\subset P$.
@@
From this point on, we will often write “vertex” to mean “vertex of $\mathrm{conv}(\mathcal{C})$” and “null” to mean “$J$-null”.
@@
Suppose there are 3 or more. Since each of them is connected to every other one by a sequence of edges, Lemma 3.5 implies that they are all proportional and thus collinear, which is impossible. ∎
@@
    Type I: one entry of w is $-1$ and the others are zero. These are denoted $(-1^{i})$ for some $i=1,\dots,r$.
@@
In our classification of superpotentials, we will assume that $\mathrm{conv}(\mathcal{W})$ has nonzero $(r-1)$-dimensional measure. This is the same hypothesis as in the main theorem of [5]. It holds, for instance, if $G$ is semisimple. In the case that the isotropy group $K$ is connected, this is not so restrictive, due to the following result.
@@
We begin by noting that there is at least one type I vector, say $(-1^{r})\in\mathcal{W}$ up to permutation of indices. Let $\textbf{w}_{1},\dots,\textbf{w}_{m}$ be the other vectors of $\mathcal{W}$. First, we note that it suffices to show that the image of $\mathrm{conv}(\mathcal{W})$ under the affine transformation which sends the last ($r$-th) coordinate to zero has nonzero measure as a subset of $\{x_{r}=0\}\cong\mathbb{R}^{r-1}$. Since $(-1^{r})$ is sent to 0 by this transformation, it suffices to show that the matrix whose rows are the components of $\textbf{w}_{1},\dots,\textbf{w}_{m}$ has rank $r-1$.
@@
Assume for sake of contradiction that there are vertices on only one side of $P$, then there are either one or two on that side. Suppose that there is just one, denoted by c. Then it is adjacent to every vertex of $\mathrm{conv}(\mathcal{C})$ in $P$, which span an $r$-dimensional subset of $P$. However, they must all lie in $\mathrm{ker}J(\textbf{c},\cdot)\cap P\subsetneq P$, which is a contradiction. To see this, use the closest point argument on the edge connecting each such vertex with c, and use the fact that c is null. Suppose that there are two, then by Lemma 3.5 they are given by c and $\lambda\textbf{c}$ for some $\lambda\in\mathbb{R}$. Then each of the vertices in $P$ must be adjacent to at least one of these, and thus be in $\mathrm{ker}J(\textbf{c},\cdot)\cap P\subsetneq P$, a contradiction for the same reason as above. ∎
@@
There is no null vertex in $P$, so since $\mathrm{conv}(\frac{1}{2}(\textbf{d}+\tilde{\mathcal{W}}))\subset\mathrm{conv}(\mathcal{C})\cap P$ and every vertex outside of $\frac{1}{2}(\textbf{d}+\tilde{\mathcal{W}})$ is null by Lemma 3.2, we have that $\mathrm{conv}(\frac{1}{2}(\textbf{d}+\tilde{\mathcal{W}}))=\mathrm{conv}(\mathcal{C})\cap P=\mathrm{conv}(\mathcal{C}\cap P)$, and that every vertex of $\mathrm{conv}(\frac{1}{2}(\textbf{d}+\tilde{\mathcal{W}}))$ is a vertex of $\mathcal{C}$. From this point on, we may follow the proof of Proposition 13 in [2], which we reproduce here for completeness. Since 0 is a vertex of $\mathrm{conv}(\tilde{\mathcal{W}})$ and there are no null vertices in $P$, there is a $\textbf{w}\in\mathcal{W}$ such that the vertices $\textbf{c}_{0}=\frac{1}{2}\textbf{d}$ and $\textbf{c}_{1}=\frac{1}{2}(\textbf{d}+\textbf{w})$ form an edge of $\mathrm{conv}(\mathcal{C})$. Then $J(\textbf{c}_{0},\textbf{c}_{0})=\frac{1}{4}$ and $J(\textbf{c}_{0},\textbf{c}_{1})=\frac{1}{4}$, and thus $J(\textbf{c}_{0},\cdot)$ is constant and nonzero on the line segment $\overline{\textbf{c}_{0}\textbf{c}_{1}}$. Consider the element of $\mathcal{C}$ on $\overline{\textbf{c}_{0}\textbf{c}_{1}}$ which is closest to $\textbf{c}_{0}$. It satisfies the unique sum condition with $\textbf{c}_{0}$, which contradicts Lemma 3.2.
It only remains to check the case where there exists a pair of vertices a and c, one on each side of $P$, such that $\overline{\textbf{a}\textbf{c}}$ is an edge of $\mathrm{conv}(\mathcal{C})$. Then either $\textbf{a}+\textbf{c}\in\textbf{d}+\tilde{\mathcal{W}}$ or a and c are proportional, as can be seen as follows.
@@
If $\textbf{x}_{a}\neq\textbf{c}$, we note that $\textbf{a},\textbf{x}_{a}$ satisfy the unique sum condition. Similarly, we may define $\textbf{x}_{c}\in\overline{\textbf{a}\textbf{c}}$ such that $\textbf{c},\textbf{x}_{c}$ satisfy the unique sum condition. Since a and c lie on different sides of $P$, we have that at least one of $\frac{1}{2}(\textbf{a}+\textbf{x}_{a})$ or $\frac{1}{2}(\textbf{c}+\textbf{x}_{c})$ is not in $\frac{1}{2}(\textbf{d}+\tilde{\mathcal{W}})\subset P$; assume without loss of generality that it is $\frac{1}{2}(\textbf{a}+\textbf{x}_{a})$. Then, by Lemma 3.2, $J(\textbf{a},\textbf{x}_{a})=0$, and by linearity $J(\textbf{a},\textbf{c})=0$, implying that a and c are proportional.
@@
Let $\textbf{f}=\frac{1}{2}(\textbf{d}+\textbf{w})$ be a non-null vertex which shares an edge with both a and c. Let $\textbf{x}_{a}$ denote the point of $\mathcal{C}$ on $\overline{\textbf{a}\textbf{f}}$ which is closest to a, and let $\textbf{x}_{c}$ denote the point of $\mathcal{C}$ on $\overline{\textbf{c}\textbf{f}}$ which is closest to c. Then we must have by Lemma 3.2 that $J(\textbf{a},\textbf{x}_{a})=0$ and $J(\textbf{c},\textbf{x}_{c})=0$. Since $J(\textbf{a},\cdot)$ and $J(\textbf{c},\cdot)$ are affine functions on the abovementioned edges and both a and c are null vertices, we must have $J(\textbf{a},\textbf{f})=J(\textbf{c},\textbf{f})=0$. This is impossible, as $J(\textbf{a},\textbf{f})+J(\textbf{c},\textbf{f})=J(\textbf{d},\frac{1}{2}(\textbf{d}+\textbf{w}))=\frac{1}{2}$. Therefore $\textbf{a}+\textbf{c}\notin\textbf{d}+\tilde{\mathcal{W}}$.
In the final case, namely that a and c are proportional, then by Lemma 3.9 there is only one vertex on each side of $P$. Furthermore, $\frac{1}{2}\textbf{d}$ is not a vertex, since otherwise there must be a non-null vertex sharing an edge with $\frac{1}{2}\textbf{d}$, which is impossible.
Let =$\ell=\{\lambda\textbf{a}\>|\>\lambda\in\mathbb{R}\}$ be the (null) line passing through a and c, intersecting $P$ at $\mathbf{l}$. Then $\textbf{l}\in\{s\geq 0\}$. Since all vertices of $\mathrm{conv}(\mathcal{C})\cap P$ except possibly l are in $\{s=-1\}$, but $\frac{1}{2}\textbf{d}\in\mathrm{conv}(\mathcal{C})\cap P$, we have that $\textbf{l}\in\{s>0\}$.
Note that $\mathrm{conv}(\mathcal{C})$ has nonzero $(r+1)$-dimensional measure, since it contains the line $\overline{\textbf{a}\textbf{c}}$ through the origin as well as $\mathrm{conv}(\tilde{\mathcal{W}})$, which is an $r$-dimensional subset of $P$. Since all vertices of $\mathrm{conv}(\mathcal{C})$ other than a and c are in $\{s=-1\}\cap P$, which is $(r-1)$-dimensional, there is a set $\mathcal{V}=\{\textbf{b}_{i}\}_{i=1}^{r}\subset\{s=-1\}\cap P$ of vertices adjacent to a such that $\dim\mathrm{conv}(\mathcal{V})=r-1$. Applying the closest point argument, we get $J(\textbf{a},\textbf{b}_{i})=0$ for each $i$. The $(r-1)$-dimensional affine subspace $\ker J(\textbf{a},\cdot)\cap P$ contains both $\mathcal{V}$ and $\textbf{l}\in\overline{\textbf{a}\textbf{c}}\cap P$, but $\textbf{l}\notin\{s=-1\}$, which is a contradiction. ∎
@@
There cannot be a non-null vertex adjacent to $\frac{1}{2}\textbf{d}$. Let a be a vertex adjacent to $\frac{1}{2}\textbf{d}$, then $J(\textbf{a},\textbf{a})=0$ and $J(\frac{1}{2}\textbf{d},\textbf{a})=\frac{1}{4}$, so that by the closest point argument and Lemma 3.2 we must have that the line segment between them is empty and that $\textbf{a}=\frac{1}{2}\textbf{d}+\textbf{w}$ for some $\textbf{w}\in\mathcal{W}$. All such vertices a, of which in all nontrivial cases $r\geq 2$ there exists more than one, therefore lie in $\{s=-2\}$. Consequently, there are at least two vertices in the half-plane $\{s<-1\}$, which are connected by edges lying entirely in $\{s<-1\}$. Thus, there is least one pair of adjacent (null) vertices in $\{s<-1\}$, which is impossible by the closest point argument and Lemma 3.2. ∎
@@
If there are 2 vertices in $\{s<-1\}$, then by the same argument they must average to an element of $\frac{1}{2}(\textbf{d}+\tilde{\mathcal{W}})$ in $\{s<-1\}$, which is impossible. Therefore $b\leq 1$. ∎
@@
More explicitly, we are saying that all vertices of $\mathrm{conv}(\frac{1}{2}(\textbf{d}+\mathcal{W}))=\mathrm{conv}(\mathcal{C})\cap\{s=-1\}$ must take the form $\frac{1}{2}(\textbf{d}+\textbf{x})$, where